# One-Versus-Others Attention: Scalable Multimodal Integration for Clinical Data

### v3

@@ Abstract
Multimodal models have become increasingly important as they surpass single-modality approaches on diverse tasks ranging from question-answering to autonomous driving. Despite the importance of multimodal learning, existing efforts focus on vision-language applications, where the number of modalities rarely exceeds four (images, text, audio, video). However, data in other domains, such as healthcare, may include many more modalities like X-rays, PET scans, MRIs, genetic screening, genomic data, and clinical notes, creating a need for both efficient and accurate data integration. Many state-of-the-art multimodal models rely on cross-attention or self-attention for effective data integration, which do not scale well for applications with more than two modalities. The complexity per layer of computing attention in either paradigm is, at best, quadratic with respect to the number of modalities, posing a computational bottleneck that impedes broad adoption. To address this, we propose a new attention mechanism, One-Versus-Others (OvO) attention, that scales linearly with the number of modalities, thus offering a significant reduction in computational complexity compared to existing multimodal attention methods. Using four clinical datasets with diverse modalities, we show that our method decreases computation costs while maintaining or increasing performance compared to popular integration techniques. Across all clinical datasets, OvO reduced the number of required floating point operations (FLOPs) by at least 91.98%, demonstrating its significant impact on efficiency and enabling wider adaptation. 111Code is available at https://anonymous.4open.science/r/OvO-13BD/.
@@ 1 Introduction
Multimodal learning has emerged as a promising approach, which enables joint learning from multiple modalities of data (e.g., text and images). Combining different modalities allows for a more comprehensive and accurate understanding of tasks such as clinical decision support (Ming et al., 2022; Golovanevsky et al., 2022; Hayat et al., 2022), image and video captioning (Yu et al., 2019; Seo et al., 2022), audio-visual speech recognition (Sterpu et al., 2018), and sentiment analysis (Poria et al., 2018). Multimodal learning has been explored through various methods in machine learning and deep learning. While feature-level integration was mostly used in more traditional machine learning algorithms, Neural Networks have allowed for the intermediate fusion of modalities at any layer and late fusion at the decision-making stage. However, these fusion paradigms lack a key component - capturing explicit interaction between modalities. For example, in detecting Alzheimer’s Disease, genetic features help reinforce and ground the clinical information and thus lead to more robust decision-making (Golovanevsky et al., 2022). Such relevant interactions can be captured through the attention mechanism. Popular multimodal models, such as LXMERT (Tan and Bansal, 2019) and ViLBERT (Lu et al., 2019), use a fusion method that captures interactions between modalities using cross-attention. On the other hand, models such as VisualBERT (Li et al., 2019) and VL-BERT (Su et al., 2019) use early fusion, where vision and language inputs are concatenated early to learn multimodal through self-attention.
@@
However, both self-attention and cross-attention grow quadratically in computational burden with the number of modalities, posing a scalability challenge. While in popular vision-language integration tasks, the number of modalities rarely exceeds four (images, text, audio, video), a significant bottleneck can arise in other domains. The healthcare domain exemplifies this issue, as a single task may involve integrating data from complex and rich sources spanning multiple modalities from radiology, pathology, genomics, genetics, and clinical data. Therefore, with the influx of many modalities, the use of cross-attention or self-attention will remain limited in the clinical domain as their computational demands escalate even further. To address this gap, we propose a new attention mechanism, One-Versus-Others (OvO) attention. OvO attention is calculated by comparing one modality against a combined representation of all other modalities (hence the name, One-Versus-Others). Our approach significantly reduces computational complexity as it grows linearly with the number of modalities (see Section 3.3). Figure 1 sketches a four-modality example to demonstrate the difference between our approach (scales linearly) and self-attention/cross-attention (scales quadratically). OvO is a general attention scheme that can be integrated into existing multimodal architectures instead of cross-attention or self-attention.
@@
We first, present a complexity analysis and demonstrate its validity through a simulated dataset. Our simulation results show scalability gains in an extreme multimodal setting (20 modalities). Next, we use four diverse clinical datasets that vary in modalities, encoder types (pre-trained and not), number of samples, and tasks to show our model’s improved scalability in different clinical multimodal settings. Our results demonstrate that our method dramatically decreases computation costs (offering at minimum a 91.98% reduction in computations), compared to self-attention and cross-attention while maintaining or even exceeding performance. Finally, we demonstrate the generalizability of our method on an additional non-clinical dataset, further underscoring OvO’s broad applicability.
@@
Overall, OvO is a novel attention scheme for multimodal integration that is domain-agnostic and scales linearly to the number of modalities. Importantly, it enables the practical application of deep learning models in healthcare, where computational efficiency and accuracy are vital for deployment.
@@ 2 Related work
Multimodal attention-based models are increasingly pivotal in clinical decision support systems, paralleling their widespread use in vision-language applications. In the medical domain, these models have shown remarkable utility in diverse scenarios, such as cancer classification (LI et al., 2021), biomarker discovery (Braman et al., 2021; Ilyin et al., 2004), prognosis prediction (Schulz et al., 2021; Silva and Rohr, 2020), and more. These applications highlight the versatility and potential of multimodal learning in handling complex and rich medical data. The attention mechanism serves as a core component in these models. Attention measures the similarity among individual representations, like word embedding vectors or, in the multimodal scenario, modality-specific embeddings. Each input embedding can assume one of three roles: (1) Query (), representing the current focus of attention when compared against other input embeddings; (2) Key (), signifying an input embedding being compared to the Query; and (3) Value (), which contributes to computing the output for the Query.
@@
Commonly, the representations from each modality in the multimodal models are passed through one of two paradigms - early fusion followed by self-attention or fusion through cross-attention. The early fusion group (e.g., Transformer-based models like Uniter (Chen et al., 2020), VisualBERT (Li et al., 2019), Vl-BERT (Su et al., 2019)) concatenates the visual embeddings and the textual embeddings as a single input, before passing the inputs through attention (see Figure 1 (a)). Given modalities and , queries (), keys (), and values () are computed from their concatenated sequence (e.g., ). The final output from a standard Transformer block is denoted by , Equation 1 shows the early fusion paradigm.
@@
The cross-attention scheme (used in Transformer-based models like ViLBERT (Lu et al., 2019), LXMERT (Tan and Bansal, 2019), ActBERT (Zhu and Yang, 2020), MulT (Tsai et al., 2019), etc.) inputs each modality into its own Transformer, the outputs of which are fed to a cross-modal Transformer (see Figure 1 (b)). For such models, the cross-modal interactions are captured in a pairwise manner through cross-attention, where queries (), keys (), and values () are computed from the modality inputs ( and ), and then the keys and values from each modality are fed to the multi-headed attention block of the other modality. The output, , is shown in Equation 2.
@@
While the early fusion and cross-attention paradigms could be extended to three modalities, seen in TriBERT (Rahman et al., 2021) and VATT (Akbari et al., 2021), these models face scalability challenges for more than three modalities. Cross-attention methods can leverage joint representations formed from cross-attention but do not scale well to larger numbers of modalities as they are computed in a pairwise fashion. Thus, if there are modalities, computing pairwise fusion between each pair will result in matrix computations. Moreover, attention is not a symmetric calculation, which means that most commonly, it is computed bi-directionally (e.g., image to text and text to image), leading to an even greater computational burden. Early fusion involves the concatenation of modalities before the Transformer layer, which similarly does not scale well with the number of modalities. Self-attention is quadratic with respect to sequence length (Vaswani et al., 2017), and since early fusion methods concatenate inputs before attention, the computational complexity will increase quadratically as the number of modalities increases (see Section 3.3). Furthermore, concatenation is not order invariant, making the ordering of modalities an important consideration, potentially requiring similar bi-directional computations as cross-attention. Our integration method, OvO, addresses the limitations mentioned above in a scalable and domain-agnostic manner.
@@ 3.1 One-Versus-Others (OvO) attention
We propose a new attention mechanism, One-Versus-Others (OvO) Attention, which grows linearly with number of modalities rather than quadratically, as would be the case for cross-attention or self-attention (see Section A). OvO computes attention between one modality at a time with respect to all other modalities. Given modality , which is an embedding obtained from a dedicated encoder, where is the number of modalities and , OvO takes in one modality and computes the dot product against all the other modalities with a weight matrix . is a learnable parameter that is shared across all modalities (see Figure 1 (c)) and can learn interactions throughout training. The similarity score function, representing the degree of alignment between the chosen modality and others, calculated for modality with respect to a set of other modalities () is shown in Equation 3. The context vector in OvO for modality , which is a combined representation of information from the other modalities, is shown in Equation 4:
We chose to sum over the “other” modalities instead of concatenation for two reasons: (1) the concatenation vector will continue to increase in length with the number of modalities, which will result in a less scalable framework; (2) concatenation is not invariant to the order of modalities, which could affect the model prediction, whereas a sum provides position invariance.
Note that for modalities (, ), the similarity score function simplifies to that of general attention (Luong et al., 2015):
@@ 3.2 Multi-headed OvO Attention
Here, is the number of attention heads, is a learnable weight matrix for the -th attention head, is a learnable weight matrix that projects the concatenated outputs of the attention heads back to the original dimension, and OvO Attention is defined in Equation 4.
@@ 3.3 Model Complexity
This section highlights the complexities of the two commonly used paradigms: early fusion followed by self-attention and pairwise cross-attention, as well as our One-Versus-Others (OvO) attention. Table 1 summarizes the complexity per layer. Let represent the number of modalities, be the feature-length of each modality (assuming equal), and be the representation dimension of the respective weight matrices. As established in (Vaswani et al., 2017), self-attention has complexity of . In the multimodal case, self-attention concatenates modalities before attention, leading to a sequence length of , influencing the quadratic term. Thus, the complexity of self-attention is . Cross-attention computes attention over all pairwise permutations of modalities: . Thus, the number of operations required by cross-attention is . When focusing on the fastest-growing terms in big notation, the final complexity per layer simplifies to . One-Versus-Others (OvO) Attention requires one attention calculation per modality, making it linear with respect to . Thus, the complexity per layer for OvO is . Appendix Section A provides step-by-step details for the complexity calculations.
| Model | Complexity Per Layer |
| Self-Attention | |
| Cross-Attention | |
| One-Versus-Others (OvO) Attention |
@@ 3.4 Illustration through simulation
To illustrate the linearity of OvO compared to the other integration paradigms, we simulated 20 artificial modalities. We consider two classes: (1) 20 random feature values that sum up to 1.0, and, (2) 20 random feature values that are each less than 0.15. These classes were created such that the correct label can only be inferred after inspecting all features. For example, 0.14 is less than 0.15, but it could also be a value that adds to 1. For more details on the simulation dataset and how the threshold was chosen, see Appendix B. Each value was then vectorized by sampling randomly around the selected number, such that each modality is a vector of size 20 rather than a single number, leading to a combined total of 400 features. Overall, the dataset contains 2,000 samples (1,000 for each class). Our constructed simulation dataset tests the scaling capabilities of our method to an extent that real-world datasets do not usually reach.
Using 2, 5, 10, 15, and 20 simulated modalities, we examine the computation cost across the three integration methods. Most notably, while self-attention and cross-attention grow quadratically with respect to the number of modalities, , (), our method scales linearly (), as shown in Figure 2.
@@ 4 Experiments
We used four diverse clinical datasets to examine our method against three standard integration techniques: concatenation with no attention (baseline), early fusion with self-attention, and pairwise cross-attention. These clinical tasks feature a range of rich modalities that, despite their high integration costs, remain essential to solve. In Section 5.1, we provide results on the Amazon Reviews dataset that show our method is applicable to fields outside the healthcare domain.
@@ 4.1 Dataset descriptions
4.1.1 The Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) dataset
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Petersen et al., 2010) database provides neuroimaging data, cognitive test scores, biomarker profiles, and genetic information for Alzheimer’s disease (AD), mild cognitive impairment (MCI), and normal patients. We use the processed data from the Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) challenge (Marinescu et al., 2019). We utilize six modalities that have the least missing information per patient: cognitive tests - neuropsychological tests administered by a clinical expert; MRI ROIs (generated from Freesurfer) - measures of brain structural integrity; FDG PET ROI averages - measure cell metabolism, where cells affected by AD show reduced metabolism; AV45 PET ROI averages - measures amyloid-beta load in the brain; demographic information (e.g., age, gender, education); and CSF biomarkers - amyloid and tau levels in the cerebrospinal fluid. The preprocessing provided by TADPOLE turned every modality into a tabular form (including imaging). After removing patients with missing modalities, we had 767 MCI patients, 493 normal patients, and 143 AD patients. Thus, this is a three-class classification task with six modalities.
4.1.2 The Cancer Genome Atlas (TCGA) dataset
TCGA is a cancer genomics program that molecularly characterized over 20,000 primary cancers and matched normal samples spanning 33 cancer types. Accurate prediction of cancer type is an important task but has primarily focused on single-modality approaches (Edge and Compton, 2010; Soh et al., 2017; Divate et al., 2022; Mohammed et al., 2021). Sun et al. (Sun et al., 2018) used multiple modalities from TCGA but did so for survival prediction rather than cancer classification. We used patients diagnosed with lung (Network et al., 2014b), stomach (Network et al., 2014a), colon (Network et al., 2012), liver (Ally et al., 2017), or kidney (working group: Baylor College of Medicine Creighton Chad J. 1 2 Morgan Margaret 1 Gunaratne Preethi H. 1 3 Wheeler David A. 1 Gibbs Richard A. 1 et al., 2013) cancer to create our dataset consisting of 5 modalities – whole-slide images, clinical data (tabular), copy number variation or CNV (tabular), gene expression (tabular), and DNA Methylation (tabular). CNV, DNA Methylation, and gene expression all describe different genomic information; full descriptions can be found in Appendix D. Clinical data includes demographics, laboratory tests, and family relationships. Imaging includes pathology slide images of tissue sampled from the tumor. In total, we had 338 colon cancer patients, 329 kidney, 301 lung, 228 liver, and 226 stomach patients after filtering described in Appendix D. This is a five-class classification task with five modalities.
4.1.3 MIMIC-IV and CXR

### v4

# One-Versus-Others Attention: Scalable Multimodal Integration for Biomedical Data

@@ Abstract
Multimodal models have become increasingly important as they surpass single-modality approaches on diverse tasks ranging from question-answering to disease diagnosis. Despite the importance of multimodal learning, existing efforts focus on vision-language applications, where the number of modalities rarely exceeds four (images, text, audio, video). However, data in healthcare domain, may include many more modalities like X-rays, PET scans, MRIs, genetic screening, genomic data, and clinical notes, creating a need for both efficient and accurate data integration. Many state-of-the-art multimodal models rely on cross-attention or self-attention for effective data integration, which do not scale well for applications with more than two modalities. The complexity per layer of computing attention in either paradigm is, at best, quadratic with respect to the number of modalities, posing a computational bottleneck that impedes broad adoption. To address this, we propose a new attention mechanism, One-Versus-Others (OvO) attention, that scales linearly with the number of modalities, thus offering a significant reduction in computational complexity compared to existing multimodal attention methods. Using three clinical datasets with multiple diverse modalities, we show that our method decreases computation costs while maintaining or increasing performance compared to popular integration techniques. Across all clinical datasets, OvO reduced the number of required floating point operations (FLOPs) by at least 91.98%, demonstrating its significant impact on efficiency and enabling multi-modal predictions in healthcare.111Code and Appendix are available at https://github.com/rsinghlab/OvO
keywords:
Multimodal learning; deep learning; attention mechanism; clinical decision support.© 2024 The Authors. Open Access chapter published by World Scientific Publishing Company and distributed under the terms of the Creative Commons Attribution Non-Commercial (CC BY-NC) 4.0 License.
@@ 1 Introduction
Multimodal learning has emerged as a promising approach, which enables joint learning from multiple modalities of data (e.g., text and images). Combining different modalities allows for a more comprehensive and accurate understanding of tasks such as clinical decision support [18, 5, 6], image and video captioning [33, 26], audio-visual speech recognition [30], and sentiment analysis [23]. Multimodal learning has been explored through various methods in machine learning and deep learning. While feature-level integration was mostly used in more traditional machine learning algorithms, Neural Networks have allowed for the intermediate fusion of modalities at any layer and late fusion at the decision-making stage. However, these fusion paradigms lack a key component - capturing explicit interaction between modalities. For example, in detecting Alzheimer’s Disease, genetic features help reinforce and ground the clinical information and thus lead to more robust decision-making [5]. Such relevant interactions can be captured through the attention mechanism. Popular multimodal models, such as LXMERT [31] and BLIP [12], use a fusion method that captures interactions between modalities using cross-attention. On the other hand, models such as VisualBERT [13] and LLaVA [14] use early fusion, where vision and language inputs are concatenated early to learn multimodal through self-attention. The clinical domain embraced these approaches, with multimodal models like MedFuseNet [27] and ARMOUR [15] using cross-attention for medical vision question answering and mortality prediction. In parallel, models such as BioViL-T [2] and MMBERT [10] employ early fusion through self-attention for disease prediction and report generation.
@@
However, both self-attention and cross-attention grow quadratically in computational burden with the number of modalities, posing a scalability challenge. While in popular vision-language integration tasks, the number of modalities rarely exceeds four (images, text, audio, video), a significant bottleneck can arise in other domains. The healthcare domain exemplifies this issue, as a single task may involve integrating data from complex and rich sources spanning multiple modalities from radiology, pathology, genomics, genetics, and clinical data. Therefore, with the influx of many modalities, the use of cross-attention or self-attention will remain limited in the clinical domain as their computational demands escalate even further. To address this gap, we propose a new attention mechanism, One-Versus-Others (OvO) attention. OvO attention is calculated by comparing one modality against a combined representation of all other modalities (hence the name, One-Versus-Others). Our approach significantly reduces computational complexity as it grows linearly with the number of modalities (see Section 3.3). Figure 1 sketches a four-modality example to demonstrate the difference between our approach (scales linearly) and self-attention/cross-attention (scales quadratically). OvO is a general attention scheme that can be integrated into existing clinical multimodal architectures instead of cross-attention or self-attention.
@@
We first, present a complexity analysis and validate it through a simulated dataset. Our simulation results show scalability gains in an extreme multimodal setting (20 modalities). Next, we use three diverse clinical datasets that vary in modalities, encoder types (pre-trained and not), number of samples, and tasks to show our model’s improved scalability in different clinical multimodal settings. Our results demonstrate that our method dramatically decreases computation costs (offering at minimum a 91.98% reduction in computations), compared to self-attention and cross-attention while maintaining or even exceeding performance.
@@
Overall, OvO is a novel attention scheme for multimodal integration that scales linearly to the number of modalities, enabling the practical application of deep learning models in healthcare, where computational efficiency and accuracy are vital for deployment.
@@ 2 Related work
Multimodal attention-based models are increasingly pivotal in clinical decision support systems, paralleling their widespread use in vision-language applications. In the medical domain, these models have shown remarkable utility in diverse scenarios, such as cancer classification [11], biomarker discovery [3, 7], prognosis prediction [25, 29], and more. These applications highlight the versatility and potential of multimodal learning in handling complex and rich medical data. The attention mechanism serves as a core component in these models. Attention measures the similarity among individual representations, like word embedding vectors or, in the multimodal scenario, modality-specific embeddings. Each input embedding can assume one of three roles: (1) Query (), representing the current focus of attention when compared against other input embeddings; (2) Key (), signifying an input embedding being compared to the Query; and (3) Value (), which contributes to computing the output for the Query.
@@
Commonly, the representations from each modality in the multimodal models are passed through one of two paradigms - early fusion followed by self-attention or fusion through cross-attention. The early fusion group (e.g., Transformer-based models like UNITER [4], VisualBERT [13], LLaVA [14], BioViL-T [2], MedViLL [20], etc.) concatenates the visual embeddings and the textual embeddings as a single input, before passing the inputs through attention (see Figure 1 (a)). Given modalities and , queries (), keys (), and values () are computed from their concatenated sequence (e.g., ). The final output from a standard Transformer block is denoted by , Equation 1 shows the early fusion paradigm.
@@
The cross-attention scheme (used in Transformer-based models like ViLBERT [16], LXMERT [31], MedFuseNet [27], MADDi [5], etc.) inputs each modality into its own Transformer, the outputs of which are fed to a cross-modal Transformer (see Figure 1 (b)). For such models, the cross-modal interactions are captured in a pairwise manner through cross-attention, where queries (), keys (), and values () are computed from the modality inputs ( and ), and then the keys and values from each modality are fed to the multi-headed attention block of the other modality. The output, , is shown in Equation 2.
@@
While the early fusion and cross-attention paradigms could be extended to three modalities, seen in TriBERT [24] and VATT [1], these models face scalability challenges for more than three modalities. Cross-attention methods can leverage joint representations formed from cross-attention but do not scale well to larger numbers of modalities as they are computed in a pairwise fashion. Thus, if there are modalities, computing pairwise fusion between each pair will result in matrix computations. Moreover, attention is not a symmetric calculation, which means that most commonly, it is computed bi-directionally (e.g., image to text and text to image), leading to an even greater computational burden. Early fusion involves the concatenation of modalities before the Transformer layer, which similarly does not scale well with the number of modalities. Self-attention is quadratic with respect to sequence length [32], and since early fusion methods concatenate inputs before attention, the computational complexity will increase quadratically as the number of modalities increases (see Section 3.3). Furthermore, concatenation is not order invariant, making the ordering of modalities an important consideration, potentially requiring similar bi-directional computations as cross-attention. Our integration method, OvO, addresses the limitations mentioned above in a scalable and domain-agnostic manner.
@@ 3.1 One-Versus-Others (OvO) attention
We propose a new attention mechanism, One-Versus-Others (OvO) Attention, which grows linearly with the number of modalities rather than quadratically, as would be the case for cross-attention or self-attention (see Section 3.3). OvO computes attention between one modality at a time with respect to all other modalities. Given modality , which is an embedding obtained from a dedicated encoder (e.g., CNN, ClinicalBERT, etc.) and where is the number of modalities, OvO takes in one modality and computes the dot product against all the other modalities with a weight matrix . is a learnable parameter that can help scale the importance of each attention calculation (see Figure 1 (c)) and can learn interactions throughout training. The modality embeddings in OvO attention function like queries, while the weighted sum of the other modalities behaves like keys and values, akin to the dot-product attention mechanisms, with the naming choice of adapted for multimodal applications. This dot product The similarity score function, representing the degree of alignment between the chosen modality and others, calculated for modality with respect to a set of other modalities () is shown in Equation 3. This produces a vector of scores which measure the relevance of with respect to the other modalities. The context vector in OvO for modality , which is a combined representation of information from the other modalities, is shown in Equation 4:
In Equation 4, the softmax is applied across the input dimension of the embeddings, normalizing across the attention scores. The result is then multiplied by the original modality embeddings to compute the final output. We chose to sum over the “other” modalities instead of concatenation because: (1) the concatenation vector will continue to increase in length with the number of modalities, which will result in a less scalable framework; (2) concatenation is not invariant to the order of modalities, which could affect the model prediction, whereas a sum provides position invariance.
Unlike cross-attention and self-attention, OvO provides a more interpretable mechanism for analyzing interactions between multiple modalities. In cross-attention, interactions are captured in isolated pairs (e.g., compared to or compared to ), limiting the ability to see how one modality integrates information from all others. Similarly, self-attention condenses modality interactions into a single operation, which can obscure explicit cross-modal interactions and make it difficult to disentangle their contributions. OvO, however, creates one attention matrix per modality, with each modality interacting with all others through the learnable weight matrix , which ensures a flexible and adaptive scaling mechanism. In future work, will be studied to better understand the relative importance of each modality’s contributions to the final prediction.
@@ 3.2 Multi-headed OvO Attention
Here, is the number of attention heads, is a learnable weight matrix for the -th attention head, is a learnable weight matrix that projects the concatenated outputs of the attention heads back to the original dimension, and Equation 4 defines OvO Attention.
@@ 3.3 Model Complexity
This section highlights the complexities of the two commonly used paradigms: early fusion followed by self-attention and pairwise cross-attention, as well as our One-Versus-Others (OvO) attention. Table 3.3 summarizes the complexity per layer. Let represent the number of modalities, be the feature-length of each modality (assuming equal), and be the representation dimension of the respective weight matrices. As established in [32], self-attention has complexity of . In the multimodal case, self-attention concatenates modalities before attention, leading to a sequence length of , influencing the quadratic term. Thus, the complexity of self-attention is . Cross-attention computes attention over all pairwise permutations of modalities: . Thus, the number of operations required by cross-attention is . When focusing on the fastest-growing terms in big notation, the final complexity per layer simplifies to . One-Versus-Others (OvO) Attention requires one attention calculation per modality, making it linear with respect to . Thus, the complexity per layer for OvO is . Appendix Section 1 provides step-by-step details for the complexity calculations.
[…]
@@ 3.4 Illustration through simulation
To illustrate the linearity of OvO compared to the other integration paradigms, we simulated 20 artificial modalities. We consider two classes: (1) 20 random feature values that sum up to 1.0, and, (2) 20 random feature values that are each less than 0.15. These classes were created such that the correct label can only be inferred after inspecting all features. For example, 0.14 is less than 0.15, but it could also be a value that adds to 1. For more details on the simulation dataset and how the threshold was chosen, see Appendix Section 2.
Each value was then vectorized by sampling randomly around the selected number, such that each modality is a vector of size 20 rather than a single number, leading to a combined total of 400 features. Overall, the dataset contains 2,000 samples (1,000 for each class). Our constructed simulation dataset tests the scaling capabilities of our method to an extent that real-world datasets do not usually reach.
We examine the computation cost across the three integration methods using 2, 5, 10, 15, and 20 simulated modalities. Most notably, while self-attention and cross-attention grow quadratically with respect to the number of modalities, , (), our method scales linearly (), as shown in Figure 2.
@@ 4 Experiments
We used three diverse clinical datasets to examine our method against three standard integration techniques: concatenation with no attention (baseline), early fusion with self-attention, and pairwise cross-attention. These clinical tasks feature a range of rich modalities that, despite their high integration costs, remain essential to solve.
@@ 4.1 Dataset descriptions
4.1.1 MIMIC-IV and CXR data